OpenStudy (anonymous):

(3x)^0 = 1 Am I right?

OpenStudy (anonymous):

So anything raised to the 0 power, always, will be 1. So yes, you are right!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Of course =)

OpenStudy (tkhunny):

The correct answer is "maybe". It depends on the value of 'x'. If x = 0, it is most assuredly NOT 1.

OpenStudy (unklerhaukus):

your right, assuming \(x\neq0\),

OpenStudy (anonymous):

Ooooo that's a good point. Anything (Not 0) raised to the 0 power is = 1. But 0 raised to the 0 power is what... undefined @tkhunny ?

OpenStudy (bibby):

0

OpenStudy (bibby):

no

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

Am I right or wrong? I'm confused.

OpenStudy (anonymous):

No your right

OpenStudy (bibby):

zero to any power is zero

OpenStudy (bibby):

I've confused myself

OpenStudy (anonymous):

Anything raised to the 0 power, not including 0, is = 1. Where (3x)^0 is 1, where x does not equal 0.

OpenStudy (unklerhaukus):

\(0^0 \) is undefined

OpenStudy (aum):

\(\large (3x)^0 = 1 \text{ for } x \ne 0\)

OpenStudy (anonymous):

I thought 0^0 was undefined.

OpenStudy (tkhunny):

0^0 is not always clear. It is referred to as "indeterminate". It is a little better than undefined. It could be undefined or it could be defined. It depends on EXACTLY the nature of how you arrived at the state of 0^0.

OpenStudy (aum):

In imprecise terms: anything raised to zero is one. zero raised to anything is zero. Then what is zero raised to zero? Is it zero or is it one?

OpenStudy (anonymous):

How about 10(mn)^0

OpenStudy (anonymous):

I don't understand this one

OpenStudy (anonymous):

As long as m or n doesn't = 0, then it is 1

OpenStudy (tkhunny):

Not quite. The requirement is that m AND n must not be zero. More plainly, m*n must not be zero.

OpenStudy (anonymous):

(9x)^0 - 9x^0 - (-9x)^0 I kind of understand this one, but I'm not sure...

OpenStudy (anonymous):

Where x does not = 0, then it's 1 -1 -1 right?

OpenStudy (anonymous):

I think the answer is -7

OpenStudy (aum):

For x not equal to 0, (9x)^0 - 9x^0 - (-9x)^0 = 1 - 9*1 - 1 = -9